# Optimizing Mixture of Block Attention

### v1

@@ 1 Introduction
Large Language Models (LLMs) (dubey2024llama3herdmodels; openai2023gpt4) are expanding into multimodal domains like video understanding (lin2023videollava; Qwen2VL) and video generation (kong2025hunyuanvideosystematicframeworklarge), requiring the ability to handle exceptionally long contexts. This vision is bottlenecked by the self-attention mechanism (vaswani2017attention), whose quadratic computational cost makes processing long sequences costly. Sparse attention (zaheer2020bigbird; guo2024blocksparse; xu2025xattention) aims to solve this by focusing computation only on important regions. Among these methods, Mixture of Block Attention (MoBA) (lu2025mobamixtureblockattention) is a promising approach where a learned router directs each query to a small subset of key-value blocks, reducing complexity to near-linear.
@@ 2 Preliminaries
This reduces complexity from $O(N^{2})$ to $O(N\cdot kB)$ when $k\ll n$. Our experiments systematically explore configurations with $B\in\{512,256,128\}$ and corresponding $k\in\{2,4,8\}$ to maintain constant sparsity of $7/8$ when $N=8192$. To encourage within-block clustering, we optionally apply a short depthwise causal convolution on keys (yang2025parallelizinglineartransformersdelta) (kernel size 3 or 5); details are provided in Appendix B.
@@ 3.2 Signal-to-Noise Ratio (SNR) Analysis
Through statistical analysis (see Appendix A), we find:
@@ 1. Tiled Top-K Selection
The top-k selection process is a primary bottleneck in the original MoBA implementation lu2025mobamixtureblockattention, which materializes a full score matrix and processes batched sequences serially. We replace this with Flash TopK (Step 1 in Figure 1), a highly optimized three-stage pipeline of fused kernels. First, a Triton kernel computes key-block centroids, producing a much smaller matrix $\mathbf{\tilde{K}}$. Second, a tiled kernel inspired by FlashAttention-2 finds the top-k key-blocks for each query by computing scores between $\mathbf{Q}$ and $\mathbf{\tilde{K}}$ without ever materializing the full score matrix to HBM, as summarized in Algorithm 3. Finally, an efficient epilogue reformats the query-centric indices into a key-block-centric varlen layout for the main attention pass. This entire pipeline is fully parallelized across batches and heads, eliminating the original performance bottleneck. The detailed breakdown of all three stages is provided in Appendix C.1.
@@ 5.1 Experimental Setup
All models are pre-trained on FineWeb-Edu (penedo2024fineweb) using AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay 0.1, and cosine learning rate schedule. We use peak LR $6\times 10^{-4}$, batch size 500K tokens. All models are trained on 100B tokens. All models use gradient clipping at 1.0, and mixed precision training with bfloat16. We evaluate on: (1) Language Modeling: WikiText2 perplexity (merity2016pointer) and 8 zero-shot tasks: OpenBookQA (OpenBookQA2018), PIQA (bisk2019piqareasoningphysicalcommonsense), HellaSwag (zellers2019hellaswagmachinereallyfinish), WinoGrande (sakaguchi_winogrande_2021), ARC-e/c (clark2018thinksolvedquestionanswering), TruthfulQA (lin2022truthfulqameasuringmodelsmimic), and LAMBADA (paperno2016lambadadatasetwordprediction). (2) Long-Context Retrieval: S-NIAH-1/2/3 from RULER (hsieh2024ruler) at 4K–64K lengths. (3) LongBench: 12 tasks (bai2024longbenchbilingualmultitaskbenchmark): single-document QA (Qasper, MField), multi-document QA (HotpotQA, 2WikiMQA, MuSiQue), summarization (GovReport, QMSum, MultiNews), few-shot (TriviaQA, SAMSum), and code (LCC, RepoBench-P).
@@ 5.2 Quality Results
Key convolution improves performance with task-specific preferences. For 340M models, kconv3 increases language modeling accuracy from 45.1% to 45.6% (Table 1), while kconv5 achieves 100% retrieval at 64K vs 85% without (Table 3). On LongBench, kconv3 reaches 15.3% (Table 5). At 1B scale, kconv3 improves LM accuracy to 52.7% (Table 2) and RULER to 68.2% (Table 4). These gains confirm convolution amplifies $\Delta\mu_{\text{eff}}$ by clustering related tokens.
@@ 7 Conclusion
Appendix A Detailed Derivation of Signal-to-Noise Ratio
@@ A.4 Retrieval Probability
Appendix B Key Convolution Design
@@ B.2 Effect on Routing
Appendix C Kernel Implementation Details
C.1 FlashMoBA Top-K Selection and Index Reformatting
@@
C.2 FlashMoBA Forward Pass Kernel
@@
C.3 FlashMoBA Backward Pass Kernel

### v2

@@ 1 Introduction
Large Language Models (LLMs) (dubey2024llama3herdmodels; openai2023gpt4) are expanding into multimodal domains like video understanding (lin2023videollava; Qwen2VL) and video generation (kong2024hunyuanvideosystematicframeworklarge), requiring the ability to handle exceptionally long contexts. This vision is bottlenecked by the self-attention mechanism (vaswani2017attention), whose quadratic computational cost makes processing long sequences costly. Sparse attention (zaheer2020bigbird; guo2024blocksparse; xu2025xattention) aims to solve this by focusing computation only on important regions. Among these methods, Mixture of Block Attention (MoBA) (lu2025mobamixtureblockattention) is a promising approach where a learned router directs each query to a small subset of key-value blocks, reducing complexity to near-linear.
@@ 2 Preliminaries
This reduces complexity from $O(N^{2})$ to $O(N\cdot kB)$ when $k\ll n$. Our experiments systematically explore configurations with $B\in\{512,256,128\}$ and corresponding $k\in\{2,4,8\}$ to maintain constant sparsity of $7/8$ when $N=8192$. To encourage within-block clustering, we optionally apply a short depthwise causal convolution on keys (yang2025parallelizinglineartransformersdelta) (kernel size 3 or 5); details are provided in Appendix C.
@@ 3.2 Signal-to-Noise Ratio (SNR) Analysis
Through statistical analysis (see Appendix B), we find:
@@ 1. Tiled Top-K Selection
The top-k selection process is a primary bottleneck in the original MoBA implementation lu2025mobamixtureblockattention, which materializes a full score matrix and processes batched sequences serially. We replace this with Flash TopK (Step 1 in Figure 1), a highly optimized three-stage pipeline of fused kernels. First, a Triton kernel computes key-block centroids, producing a much smaller matrix $\mathbf{\tilde{K}}$. Second, a tiled kernel inspired by FlashAttention-2 finds the top-k key-blocks for each query by computing scores between $\mathbf{Q}$ and $\mathbf{\tilde{K}}$ without ever materializing the full score matrix to HBM, as summarized in Algorithm 3. Finally, an efficient epilogue reformats the query-centric indices into a key-block-centric varlen layout for the main attention pass. This entire pipeline is fully parallelized across batches and heads, eliminating the original performance bottleneck. The detailed breakdown of all three stages is provided in Appendix D.1.
@@ 5.1 Experimental Setup
All models are pre-trained on FineWeb-Edu (penedo2024fineweb) using AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay 0.1, and cosine learning rate schedule. We use peak LR $6\times 10^{-4}$, batch size 500K tokens. All models are trained on 100B tokens. All models use gradient clipping at 1.0, and mixed precision training with bfloat16. We evaluate on: (1) Language Modeling: WikiText2 perplexity (merity2016pointer) and 8 zero-shot tasks: OpenBookQA (OpenBookQA2018), PIQA (bisk2019piqareasoningphysicalcommonsense), HellaSwag (zellers2019hellaswagmachinereallyfinish), WinoGrande (sakaguchi_winogrande_2021), ARC-e/c (clark2018thinksolvedquestionanswering), TruthfulQA (lin2022truthfulqameasuringmodelsmimic), and LAMBADA (paperno2016lambadadatasetwordprediction). (2) Long-Context Retrieval: S-NIAH-1/2/3 from RULER (hsieh2024ruler) at 4K–64K lengths. (3) LongBench: 12 tasks (bai2023longbenchbilingualmultitaskbenchmark): single-document QA (Qasper, MField), multi-document QA (HotpotQA, 2WikiMQA, MuSiQue), summarization (GovReport, QMSum, MultiNews), few-shot (TriviaQA, SAMSum), and code (LCC, RepoBench-P).
@@ 5.2 Quality Results
Key convolution improves performance with task-specific preferences. For 340M models, kconv3 increases commonsense reasoning from 45.1% to 45.6% (Table 1), while kconv5 achieves 100% retrieval at 64K vs 85% without (Table 3). On LongBench, kconv3 reaches 15.3% (Table 5). At 1B scale, kconv3 improves LM accuracy to 52.7% (Table 2) and RULER to 68.2% (Table 4). These gains confirm convolution amplifies $\Delta\mu_{\text{eff}}$ by clustering related tokens.
@@ 7 Conclusion
Reproducibility Statement.
To ensure reproducibility of our work, we provide comprehensive details throughout the paper and appendices. Our theoretical contributions include complete proofs: the SNR formula derivation is presented in Section 3 with full mathematical details in Appendix B. All model architectures and hyperparameters are specified in Section 5, including the hybrid SWA-MoBA design, hidden dimensions, and head configurations. Training details are fully described: we use publicly available FineWeb-Edu dataset (penedo2024fineweb) with all experiments using 100B tokens. Our evaluation uses standard benchmarks with exact task specifications: RULER (hsieh2024ruler) and LongBench (bai2023longbenchbilingualmultitaskbenchmark). The FlashMoBA kernel implementation is detailed in Section 4 with algorithmic pseudocode, and complete implementation details appear in Appendix B. We will release our code including model implementations, FlashMoBA CUDA kernels, and evaluation scripts upon publication. All experiments were conducted on 8× H100 80GB GPUs with PyTorch and FlashAttention-2 (dao2023flashattention2), enabling exact reproduction of our results.
Appendix A LLM Usage Statement
We acknowledge the use of Large Language Models (specifically Claude and GPT-5) in the preparation of this manuscript. The LLMs were used exclusively as writing assistants to:
  • Polish and refine the language for clarity and conciseness
  • Improve grammar and sentence structure
  • Suggest alternative phrasings for technical descriptions
  • Help organize and structure sections for better flow
Appendix B Detailed Derivation of Signal-to-Noise Ratio
@@ B.4 Retrieval Probability
Appendix C Key Convolution Design
@@ C.2 Effect on Routing
Appendix D Kernel Implementation Details
D.1 FlashMoBA Top-K Selection and Index Reformatting
@@
D.2 FlashMoBA Forward Pass Kernel
@@
D.3 FlashMoBA Backward Pass Kernel